OpenStudy (anonymous):

3/4 - 1/2 * 7/8 simplest form

OpenStudy (lizz123):

Multiply first straight across

OpenStudy (anonymous):

@stylestar5678

OpenStudy (lizz123):

\[\frac{ 1 }{ 2 }\times \frac{ 7 }{ 8 }\]

OpenStudy (lizz123):

It is equals what

OpenStudy (anonymous):

7/16

OpenStudy (lizz123):

correct

OpenStudy (lizz123):

\[\frac{ 7 }{ 16 } - \frac{ 3 }{ 4 }\] change the denominators

OpenStudy (lizz123):